Suppose a savings and loan pays a nominal rate of

3.9​%

on savings deposits. Find the effective annual yield if interest is compounded quarterly.
Answer #1

Effective annual yield (in %):

= (1.039)^4 - 1

= 1.16536 - 1

= .1654 (Rounded Off)

= 16.54%

Effective Annual Rate of Interest:

Formula: i = (1 + r/n)^n - 1

where r = nominal rate of interest

n = no.of periods compunding

= (1 + .039/4)^4 - 1

= ( 1.00975)^4 - 1

= 0.03957

= 3.957%